In a business,Romeo and Juliet gained some amount in a certain ratio.Juliet and frank received the profits in the same ratio as Romeo and Juliet.if Romeo received $250 and frank received $360,what is the amount received by Juliet?
Answer by Theo(13342)

let r = the amount that romeo gets.
let j = the amount that juliet gets.
let f = the amount that frank gets.

the ratio between romeo and juliet is r/j.

the ratio between juliet and frank is j/f.

these ratios are equal, so you get r/j = j/f.

if you want to find what juliet gets, you need to solve this equation for j.

multiply both sides of the equation by j and multiply both sides of the equation by f to get r * f = j * j

simplify to get r * f = j^2

solve for j to get j = plus or minus sqrt(r * f).

minus doesn't makes sense, so stick to j = plus sqrt (r * f).

since r = 250 and f = 360, this equation becomes j = sqrt (250 * 360).

solve for j to get j = 300.

that's your solution.

you now have the original ratios as 250/300 = 300/360.

simplify these fractions to get 5/6 = 5/6.

the ratios are the same.

the solution looks good.

the solution is that juliet gets 300.